Given the functions f(x)= sqrt of(x+8)and g(x)= 7x - 4, find f(g(4)).
If a decimal is needed, maintain 2 decimal places accuracy.

Answer by edjones(8007)

f(x)= sqrt of(x+8)and g(x)= 7x - 4
g(4)=28-4=24
f(24)=sqrt(24+8)=sqrt(32)=sqrt(16*2)=4sqrt(2)
Ed